[Brahms-l] Shift report 20050129 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Jan 29 2005 - 15:59:28 EST
 Had beam for 3 hours at the beginning of shift. 
 After the beam was dumped, there was an acces when Kris tried 
to fix the problem with the mainframe MRS1440. He replaced the 
card for the TRMRS, but it's still not working. After the acces, 
the machine people tried to give us beam, but they 
had 2 Yellow QLI.

 Data runs:
 FS 8deg@A1723
   Run       Tr. 2
   12665     100k
   12666     57k

 Took also pedestal runs (12667 and 12668).


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 29 15:59:48 2005

This archive was generated by hypermail 2.1.8 : Sat Jan 29 2005 - 16:00:10 EST